Ford's North American operation import automotive parts from Mazda (Japan) every month, year after year. They agree that all purchases made by Ford will be made in Japanese Yen at the spot rate as long as the FX rate is between $.0080/yen and $.0087/yen. If, however, the FX rate falls outside this range, Ford and Mazda will share the difference equally. Ford has an account payable of Yen 25,000,000 for the month of March.

1. The spot rate at the payment date is .0095/yen. According to the contract, how much US$ will Ford need to prepare in order to pay Mazda?

2. If there is no contract between Ford and Mazda Japan, how much US$ Ford will need to prepare?
Answer #1

Given, Spot rate at the payment date = $ 0.0095 / yen and as per contract it is out of range of FX rate ( $ 0.0080 and $0.0087)

The difference of Spot rate outside of contract range = $ 0.0095 - $.0087 = $0.0008 / yen .

Since as per contract it is shared equally if outside of agreed FX range, so each share = $ 0.0008/2 = $ 0.0004 / yen

Ford has an account payable of yen 25,000,000 = 25 x 106 yen

1. So, as per contract Ford need to prepare = $ 0.0004 x 25 x 106 + 0.0087 x 25 x 106   = $ 10,000 + 217500= $ 227500

2. If there is no contract then Ford need to prepare = $ 0.0095 x 25 x 106 = $ 237500